OpenStudy (anonymous):

2(x-4)-(x+2)=(x+2)(x-4)...I need to know if this has a solution or not and if yes what it possibly is..i had an assignment and got a bunch wrong, and I am going through to see what I did wrong

OpenStudy (anonymous):

This has no solution.

OpenStudy (anonymous):

Why is it saying it is wrong, cause I got the same thing...this is confusing

OpenStudy (anonymous):

If you bring all x's onto the left side, expand, and simplify, you get 5*x-12-x^2=0 Solving for x yields imaginary solutions.

OpenStudy (anonymous):

Stop, do not pass go, the expression does have solutions. Made arithmetic error.

OpenStudy (anonymous):

x -10 = x^2 -2x -8 x^2 -3x +2 =0

OpenStudy (anonymous):

(x-2)(x-1)=0 x=1,2